OpenStudy (tacobell518):

could some one solve this 5x + 1 = 6x - 3

OpenStudy (samantha2234):

Is their any parentheses?

OpenStudy (tacobell518):

not that I know of it ends up beig one number

OpenStudy (samantha2234):

I can't really find a solution... can you check again?

OpenStudy (tacobell518):

let me give you the whole thing

OpenStudy (samantha2234):

ok

OpenStudy (tacobell518):

Quadrilateral QRST is a rectangle. RT = 5x + 1 and QS = 6x – 3. Find QS.      8 21 42 4

satellite73 (satellite73):

\[5x + 1 = 6x - 3\] takes exacty two steps to solve

OpenStudy (samantha2234):

its 4

satellite73 (satellite73):

a) subtract \(5x\) from both sides b) add \(3\) to both sides that is all

OpenStudy (tacobell518):

thanks

OpenStudy (samantha2234):

RT=QS 5x+1=6x-3 4=x QS=6*4-3=21

OpenStudy (tacobell518):

so its 21

OpenStudy (samantha2234):

ya

OpenStudy (tacobell518):

thank you so much

OpenStudy (samantha2234):

^.^ good luck, and next time you can search google as well, you'll get your answer faster. No problem :)